Calculate nva at fc
Sales =1000
Output used for self cons.by owner =100
Decrease in stock =1500
Purchase of material=500
Electricity charges=30
Depreciation=60
Ecxise duty=20
Income =10
1
Expert's answer
2019-09-11T10:37:39-0400

GVA =1000+100-1500-500-30-60-20 = 1100-2110=-1010

NVA=GVA - depreciation = -1010-60=-950

negative NVA means thar firm run business unsucessful